OpenStudy (anonymous):

Solve by the Substitution Method: -x+2y+3z=1 z=-2 2x+y+3z=15

OpenStudy (anonymous):

Help?:>

OpenStudy (anonymous):

-x+2y-6=1 ..............> 1 therefore: -x+2y=7 ....so.... x=2y-7 2x+y-6=15................>2 therefore:2x+y=21 therefore bu Sub. in 2 2(2y-7)+y=21 .......so...... 5y-7=21 therefore: 5y=35 y=7 therefore: x=7

OpenStudy (anonymous):

Thank you so much. (:

OpenStudy (anonymous):

wlcm :)